8-pack control system working group Meeting Notes

26 August 2003


In attendance: R. Chestnut, R. Hall, J. Nelson, J. Rock

Note taker/typer: J. Nelson




*trickle down of SLED and phase distrib numbers

Ron reported ETA is ~3 weeks and that Kukhee is working on it.

(see minutes from 8/19/03 for spec)

*AB/EPICS check out plans

††††† Judy and Janice got a list, but itís not clear how helpful it is.Waiting for a bit more help from Serge, hopefully in the next few days.

*TR07 not connected

††††† Ron looked after the meeting and reseated a network connection then it worked.Heís going to ask Terri about it.



*AES matlab routines (mca and mar)

††††† Judy and Bob will help Janice get buttons on an epics panel which will call the executables she has for looking at AES data

*purple status from AB

††††† (no new news)

Janice is supposed to harass Larsen (?) about Casselís satellite crate of AB modules that makes Alan Hillís DCM code not work.

††††† Judy created a PV and itís being cmlogged to show when nlcta plc statuses change to purple.

*the saturate numbers donít go to Ė1 before they go to their values

††††† Ron will check with Kukhee

*save sensor number instead of module and channel information for the largest rms

††††† This has been fixed.

*cmlog not so useful after mod 1 fire

††††† (this is ongoing) Judy and Ron are working on code to watch cmlog for the logging of the heartbeat PV as well as an ALH branch for NLCTA processes: cmlog, alh, channel watcher, etc.

*new DTIZ module (MacNairís)

††††† Ron has info from Dave and Linda about how to read out the modules.Kukhee will write the driver.ETA ~3 weeks.Janice will ask Marc for a charge number for the work

*cmlog tutorial needed?

††††† Janice will work with Ron MacKenzie if a mini cmlog class is needed for Janice &/or nlcta operations group



(still not done) Janice committed to looking for missing 8pack variables and will double check with Steve Smith to make sure they donít want to keep any of the 8pack SIS waveforms from a while ago


From before:

SWE is ready to test processing loop with real signals.Sounds like there might not be time to test the loop until mid-October.


Steve reported that the LLRF calibration will be done in matlab.Will need SWE support to add buttons to the EPICs panels to call these.



-Once the RMS calcs for nlcta AES waveforms is set, weíll migrate the code for the 8-pack AES

-(no change) We deferred the following: longer term issue: what to do if reading out an SIS when a fault occurs?



DAC/ADC Issues

All of these issues are for sooner see the comment above: (SEE COMMENTS ABOVE)

- The trickle down of the SLED variables.This was specified at a previous meeting (numbers on SLED panel need to trickle down to I/Q and A/P PVs).For processing, we might want this feature functional sooner rather than later.

-The trickle down of the variables on the power distribution panel as well.There are 4 PVs there called Klystron Sled-in phases.These need to trickle down to the phase DACs as well.This phase is to be added to the channel phase angle (IQ calib panel) and to the timestep phase before going out to the DAC.



-The SW for the deadman is done.Is the hardware done?


Peak Power Meter (PPM)

-Janice reported that both PPMs are hooked up to labview, though itís not running all the time during commissioning and load processing.

-Tonee has yet to write the labview code to retrieve the scope waveforms, but is working on it.

-Only the PPA PARMs were being archived.Janice gave the new list to Judy/Bob.Once the waveform retrieval is working, weíll add those PVs to the archive list.

-Janice will teach llrf how to get at archived data and panel.



-(no change) On going: Need to check out readback with real signals, check archiving of variables (pwr, pmin, pmax, etc), and check the plots.


Archiving of SISs

Need to ask Joe about the following:

-(no change) Later SWE will work with Joe on his other two requests: archive every few minutes and when a change occurs in a selected set of values in EPICS (a pseudo-trip if you will)